The 2014 balance sheet of Maria’s Tennis Shop, Inc., showed $479024 in the common stock account and $3.9 million in the additional paid-in surplus account. The 2015 balance sheet showed $481970 and $3.9 million in the same two accounts, respectively. If the company paid out $348466 in cash dividends during 2015, what was the cash flow to stockholders for the year
Answer #1

Total equity=Common stock+Additional paid in surplus

Hence Total equity for 2014=(479024+3,900,000)=$4,379,024

Total equity for 2015=(481970+3,900,000)=$4,381,970

Hence net new equity=(4,381,970-4,379,024)=$2946

Hence cash flow to stockholders=Cash dividend-net new equity

=(348466-2946)

=$345520.